# FonTS: Text Rendering with Typography and Style Controls

### v1

@@ Abstract
Visual text images are prevalent in various applications, requiring careful font selection and typographic choices. Recent advances in Diffusion Transformer (DiT)-based text-to-image (T2I) models show promise in automating these processes. However, these methods still face challenges such as inconsistent fonts, style variation, and limited fine-grained control, particularly at the word level. This paper proposes a two-stage DiT-based pipeline to address these issues by enhancing controllability over typography and style in text rendering. We introduce Typography Control (TC) finetuning, an efficient parameter fine-tuning method, and enclosing typography control tokens (ETC-tokens), which enable precise word-level application of typographic features. To further enhance style control, we present a Style Control Adapter (SCA) that injects style information through image inputs independent of text prompts. Through comprehensive experiments, we demonstrate the effectiveness of our approach in achieving superior word-level typographic control, font consistency, and style consistency in Basic and Artistic Text Rendering (BTR and ATR) tasks. Our results mark a significant advancement in the precision and adaptability of T2I models, presenting new possibilities for creative applications and design-oriented tasks.
@@ 1 Introduction
Previous efforts have mainly focused on improving control over the accuracy of scene text rendering [44, 38, 8, 9]. With the development of DiT-based T2I models, e.g. SD3 [12] and Flux .1 [1], the accuracy of text rendering has seen significant improvements. Beyond content accuracy, Glyph-ByT5 [20] introduced a new text encoder pre-trained using contrastive learning, enabling various font types in text rendering. Textdiffuser-2 [9] trained both two language models and the whole diffusion model to acquire layout planing capabilities. While these methods [20, 9] have implemented control at the paragraph-level (more than 10 words), they have not yet realized word-level control. Moreover, prior text rendering methods often overlook the artistic aspects of text rendering [9]. Recent DiT models [12, 1] have demonstrated promising capabilities in artistic text rendering, yet they face challenges such as semantic confusion and style inconsistency.
To expand the boundaries of existing methods and provide enhanced control on image generation, this paper identifies three essential requirements that text rendering methods: 1) control of fonts and word-level attributes in Basic Text Rendering (BTR); 2) consistency in style control in Artistic Text Rendering (ATR); 3) preservation of Scene Text Rendering (STR) capabilities without negative impact.
To this end, we propose a two-stage DiT-based pipeline for text rendering with typography and style controls. To achieve typography control, we introduce Typography Control (TC)-finetuning, an efficient parameter fine-tuning method, alongside enclosing typography control tokens (ETC-tokens) and paired typography control dataset (TC-dataset). Our results demonstrate that the model not only learns typographic elements but also applies specific typographic features at precise word locations. To ensure content accuracy while enhancing style control, we introduces a style control adapter (SCA) that injects style information via images, independent of the text content in the prompt. For training the SCA, we collected a dataset of approximately 600k image-text pairs with high aesthetic scores.
We validate the effectiveness of the proposed methods. First, we demonstrate that the learned ETC-tokens can generate text images with the desired word-level typographic attributes, through GPT-4o [26] and manual verification. Next, we assess font consistency in BTR and style consistency in ATR by conducting user studies. These evaluations show that our method outperforms various baseline methods in terms of font consistency and word-level controllability for BTR, as well as style consistency for ATR.
@@
  • We present a DiT-based pipeline that enhances the controllability of typography and style in text rendering while preserving the capabilities of scene text rendering.
  • We propose a parameter-efficient technique that enables DiT-based T2I models to learn and apply word-level typographic attributes. The use of ETC-tokens allows model to learn specific attributes and the precise locations.
  • Comprehensive experiments validate the effectiveness of our method via superior performance in typography and style control compared to existing baselines. This is the first work to achieve word-level control in text rendering and provides a valuable reference for future research.
@@ 2 Related Work
Scene Text Rendering. Despite progress in diffusion models [31, 27], high-quality scene text rendering remains a challenge. To address this, one line of research [44, 9, 8, 37] focuses on explicitly controlling the position and content of the text being rendered. For instance, methods such as GlyphControl [44] and AnyText [37] leverage glyph images containing multiple text lines as priors to guide diffusion models to produce accurate accurate text, relying on ControlNet [46]. Similarly, TextDiffuser [9, 8] employs character-level segmentation masks as control conditions in scene text rendering. Another line of works [20, 21] fine-tune the character-aware ByT5 encoder [19] using paired glyph-text datasets, improving the ability to render accurate text in various real-world images.
While these methods improve text content accuracy in scene text rendering, they have led to a trade-off in model performance on stylization, limiting text style diversity and complicating the rendering of artistic text images [8].
@@
Artistic Text Rendering. Early research focused on font creation by transferring textures from existing characters, employing stroke-based methods [5, 33] and patch-based techniques [40, 41, 42]. The advent of Generative Adversarial Networks (GANs) improved the synthesis of new glyphs from limited examples and the separation of style and content in font design, enhancing flexibility and realism [2, 15, 14, 43, 22, 35]. Innovations with diffusion models [36, 39, 25] have enabled diverse text image stylization and semantic typography, resulting in visually appealing designs that retain readability. However, despite recent DiT models [12, 1] showing quite promise in artistic text rendering, they still struggle with semantic confusion and style inconsistency.
Our work differs from previous research: 1) Rather than stylizing an existing text glyph image, we generate artistic text images based on the text rendering capabilities of the DiT model (Flux [1]). 2) We trained style adapters on large-scale image-text pairs to handle various style conditions, including artistic text images and arbitrary style images.
Controllable Text Rendering and Image Generation. Controllable text rendering. Glyph-ByT5 [20] introduces a text encoder trained through contrastive learning, enabling paragraph-level (more than ten words) control over font types. Textdiffuser-2 [9] trains two language models with the full U-Net backbone to achieve layout planning capabilities for text rendering. While these methods have successfully enabled paragraph-level control, challenges remain in achieving fine-grained word-level control.
@@
Text-based controllable generation. Text-based controllable generation methods customize image outputs by fine-tuning diffusion models using user-provided examples. [13, 32, 16]. These approaches introduce modifier tokens to guide the generation process. ColorPeel [6] further enhances this by constructing datasets of color-shape pairs and corresponding color-prompt tokens, enabling the generation of images with precise target colors. Unlike methods that use a single modifier token to represent an object [16, 13] or a color [6], we propose using enclosing modifier tokens. Each enclosing token consists of a starting and an ending token to represent typographic attributes. This approach allows the model to not only learn the attribute but also understand its precise application location, such as a specific word.
Image-based controllable generation. Methods for image-based control, such as UniControl [28] retrain T2I models from scratch, are computationally expensive [46]. A efficient alternative introduces trainable modules to existing architectures [24, 45, 46, 47]. These adapter-based methods enable structural control [46, 24] and style control from images [45, 7]. Most of them have been implemented on U-Net backbones, with limited exploration of DiT-based T2I models. Additionally, controllable generation based on multi-modal conditions remains underexplored.
Our work extends existing controllable generation methods to DiT-based T2I models, enabling fine-grained, word-level typographic control. We also integrate both text-based control and image-based control into the generation process, broadening the applicability in generated contents.
@@ 3 Approach
Our proposed parameter-efficient fine-tuning method with enclosing typography control tokens (ETC-tokens), shown in Figure 2 (a). In addition, we detail the training of style control adapters for DiT blocks of the transformer backbone, depicted in Figure 2 (b).
@@ 3.1 Typography Control Learning
Typography Control Fine-tuning. While previous studies have shown that fine-tuning specific U-Net components can generate specific objects and colors through learned prompts [6, 16, 13], approaches are not suitable for our task due to architectural differences between DiT and U-Net. Following [16, 17], we analyzed parameter changes in the fine-tuned transformer backbone on the target dataset for 100k steps using the loss ${L_{CFM}}$ in Eq. 2. The change in parameters for layer $l$ is calculated as $\Delta_{l}=||\theta_{l}^{\prime}-\theta_{l}||/||\theta_{l}||,$, where $\theta_{l}^{\prime}$ and $\theta_{l}$ are the fine-tuned and pretrained model parameters, respectively. The total change across all layers is: $\Delta_{sum}=\sum_{l=0}^{n}mean(\Delta_{l})$. These parameters are derived from two types of layers: (1) MM-DiT blocks (merging text and image embeddings), and (2) Single-DiT blocks (processing merged embeddings from MM-DiT). In MM-DiT blocks, parameters are divided into three components: joint text attention (Txt-Attn), joint image attention (Img-Attn), and additional modules like multi-layer perception (MLP) and modulation blocks. Figure 3 shows that MM-DiTs have approximately double the weight change of Single-DiTs, with the Txt-Attn component showing nearly twice the change of other MM-DiT elements, despite it is only 5% of the total parameters in the transformer backbone.
@@
[…]
To fill the gap in high-quality datasets that combine text with typographic attributes, we created the TC-Dataset using typography control rendering (TC-Render). The pipeline leverages HTML rendering to produce images featuring typographic attributes like fonts and word-level styles such as bold, italic, and underline, as shown in Figure 4. Details of the dataset are available in the supplementary.
@@ 3.2 Style Control Adapters
$\vspace{-1.5mm}\begin{split}{z}^{\prime\prime}=\text{Attention}({Q}^{\prime},{K}^{\prime},{V}^{\prime})=\text{Softmax}(\frac{{Q^{\prime}}{K^{\prime}}^{\top}}{\sqrt{d}}){V}^{\prime},\\ \end{split}$ (4)
@@
where ${Q}={z_{c}}{W}_{q},{K}={z_{c}}{W}_{k},{V}={z_{c}}{W}_{v},{Q}^{\prime}={z_{t}}{W}_{q},{K}^{\prime}={c}_{img}{W}^{\prime}_{k},{V}^{\prime}={c}_{img}{W}^{\prime}_{v}$ , and $\lambda$ represents the influence of $c_{img}$. We freeze the base model, only the ${W}^{\prime}_{k}$ and ${W}^{\prime}_{v}$ are trainable.
Methods / $\frac{\textbf{Task}}{\textbf{Metrics}}$ Basic Text Rendering Artistic Text Rendering
@@
Glyph-ByT5 [20] 96.36 - 32.73 - -
TextDiffuser-2 [9] 42.86 - 1.81 - -
SD3-medium [12] 48.05 - 0.91 24.16 13.64
Flux.1-dev [1] 66.49 - 0.91 48.71 17.42
Ours 82.85 55.00 63.64 61.78 74.43
Table 1: Comparison of quantitative results and user study outcomes. The best results are indicated in bold formats.
@@
Flux.1-dev [1] 24.17 29.93
@@
Figure 6: Qualitative results on the artistic text rendering task compared with Midjourney [23], Flux [1] and SD3 [12].

### v2

@@ Abstract
Visual text rendering are widespread in various real-world applications, requiring careful font selection and typographic choices. Recent progress in diffusion transformer (DiT)-based text-to-image (T2I) models show promise in automating these processes. However, these methods still encounter challenges like inconsistent fonts, style variation, and limited fine-grained control, particularly at the word-level. This paper proposes a two-stage DiT-based pipeline to address these problems by enhancing controllability over typography and style in text rendering. We introduce typography control fine-tuning (TC-FT), an parameter-efficient fine-tuning method (on $5\%$ key parameters) with enclosing typography control tokens (ETC-tokens), which enables precise word-level application of typographic features. To further address style inconsistency in text rendering, we propose a text-agnostic style control adapter (SCA) that prevents content leakage while enhancing style consistency. To implement TC-FT and SCA effectively, we incorporated HTML-render into the data synthesis pipeline and proposed the first word-level controllable dataset. Through comprehensive experiments, we demonstrate the effectiveness of our approach in achieving superior word-level typographic control, font consistency, and style consistency in text rendering tasks. The datasets and models will be available for academic use.
@@ 1 Introduction
Previous efforts have mainly focused on improving control over the content accuracy of scene text rendering [48, 42, 9, 10]. With the development of DiT-based T2I models, e.g. SD3 [15] and Flux .1 [1], the accuracy of text content has seen significant improvements. Beyond content accuracy, Glyph-ByT5 [23] introduced a new text encoder through contrastive learning, enabling various font types of text. Textdiffuser-2 [10] trained both two language models and the whole diffusion model to acquire layout planing capabilities. While these methods [23, 10] have implemented control at the paragraph-level, no methods have yet realized word-level control. Moreover, prior methods often overlook the artistic aspects of text  [10]. Recent DiT models [15, 1] have demonstrated promising capabilities in artistic text rendering, yet they face challenges like semantic confusion and style inconsistency.
To expand the boundaries of existing methods (summarized in Table 1), this paper identifies three essential requirements of text rendering methods: 1) control of fonts and word-level attributes in Basic Text Rendering (BTR); 2) consistency in style control in Artistic Text Rendering (ATR); 3) preservation of Scene Text Rendering (STR) capabilities without negative impact.
To this end, we propose a two-stage DiT-based pipeline for text rendering with typography and style controls. For typography control, we introduce Typography Control (TC)-finetuning, an parameter-efficient fine-tuning method, alongside enclosing typography control tokens (ETC-tokens). By introducing HTML-render to ingeniously design the data synthesis pipeline, we propose the first word-level typography control dataset (TC-dataset). Our findings show that the model not only learns typographic elements but also applies specific typographic features at precise word locations. For style control, we introduces a style control adapter (SCA) that injects style information without compromising the accuracy of the text. The training of SCA is also a two-stage process, each stage using a different dataset. In total, these datasets consist of approximately 600k image-text pairs with high aesthetic scores.
We validate the effectiveness of the proposed methods. First, we demonstrate that the learned ETC-tokens can generate text images with the desired word-level typographic attributes, through GPT-4o and manual verification. Next, we assess font consistency in BTR and style consistency in ATR by user studies and quantitative metrics. These evaluations show that our method outperforms various baselines in terms of font consistency and word-level controllability for BTR, and style consistency for ATR.
@@
  • We are the first to address the challenge of word-level control in text rendering, via introducing a two-stage DiT-based pipeline that ensures consistency in font and style while preserving scene text rendering capabilities.
  • We propose a parameter-efficient fine-tuning technique that enables DiT-based T2I models to achieve precise control over local visual details, such as word-level typographic attributes. To address style inconsistency, we design a text-agnostic SCA that prevents content leakage while enhancing style consistency.
  • We introduce the first word-level controllable dataset. By leveraging ETC-tokens, we enables precise learning of typographic attributes and their specific locations.
  • Our approach outperforms existing baselines, demonstrating superior performance in text rendering while achieving enhanced control over typography and style.
@@ 2 Related Work
[…]
Ds-Fusion [ICCV 23]
Font-Studio [ECCV 24]
AnyText [ICLR 24]
Textdiffusers-2 [ECCV 24]
Glyph-ByT5 [ECCV 24]
SD3 / Flux [ICML 24]
Ours ✓+C ✓+C ✓+C
Table 1: Differences with existing methods, C means controls.
@@
Scene Text Rendering. Despite progress in diffusion models [35, 31], high-quality scene text rendering remains a challenge. To address this, one line of research [48, 10, 9, 41] focuses on explicitly controlling the position and content of the text being rendered, relying on ControlNet [51]. Another line of works [23, 24] fine-tune the character-aware ByT5 text encoder [22] using paired glyph-text datasets, improving the ability to render accurate text in images.
@@
Artistic Text Rendering. Early research focused on font creation by transferring textures from existing characters, employing stroke-based methods [6], patch-based techniques [44, 45, 46], and GAN-based [3, 18, 17, 47, 25] methods. Innovations with diffusion models [38, 43, 28] have enabled diverse text image stylization and semantic typography, resulting in visually appealing designs that retain readability. However, despite recent DiT models [15, 1] showing quite promise in artistic text rendering, they still struggle with semantic confusion and style inconsistency.
@@
Controllable Image Generation. Text-based controllable methods [16, 36, 19] customize image outputs by fine-tuning diffusion models using user-provided examples. These approaches introduce modifier tokens to guide the generation process. ColorPeel [7] further enhances this by constructing datasets of color-shape pairs, to generate images with target colors. Image-based controllable methods, such as UniControl [32] retrain T2I models from scratch, are computationally expensive [51]. A efficient alternative introduces trainable modules to existing architectures [27, 49, 51]. These adapters enable structural  [51, 27] and style control from images [49, 8].
Most prior approaches are implemented on U-Net with a single CLIP text encoder. There has been relatively limited exploration of DiT-based T2I models that can incorporate multiple text encoders. Moreover, the area of controllable generation under multi-modal conditions has not been well-explored. Our work extends them to DiT-based models, enabling word-level typographic control, a more fine-grained form of control than previously achieved. Additionally, we seamlessly integrate both text-based and image-based controls to expand the range of real-world applications.
@@ 3 Approach
Our proposed pipeline trains distinct components for different objectives to achieve uniquely balance between the content accuracy and stylization. The proposed parameter-efficient fine-tuning method with enclosing typography control tokens (ETC-tokens), shown in Figure 2 (a), provides word-level controls under resource constraints. Meanwhile, style control adapters training (in Figure 2(b)) overcomes the content leakage in style control.
@@ 3.1 Typography Control Learning
Typography Control Fine-tuning. Previous studies have shown that fine-tuning certain U-Net components can generate specific objects and colors through learned prompts (modifier tokens) within single CLIP text encoder  [7, 19, 16]. However, these methods are not applicable to our pipeline. The reason for this lies in the the architectural disparities between DiT and U-Net, and also due to differences between T5 and CLIP. Following [19, 20], we analyzed parameter changes in the fine-tuned transformer backbone on the target dataset for 100k steps using the loss ${L_{CFM}}$ in Eq. 2. The change in parameters for layer $l$ is calculated as $\Delta_{l}=||\theta_{l}^{\prime}-\theta_{l}||/||\theta_{l}||,$, where $\theta_{l}^{\prime}$ and $\theta_{l}$ are the fine-tuned and pretrained model parameters, respectively. The total change across all layers is: $\Delta_{sum}=\sum_{l=0}^{n}mean(\Delta_{l})$. These parameters are derived from two types of layers: (1) MM-DiT blocks (merging text and image embeddings), and (2) Single-DiT blocks (processing merged embeddings from MM-DiT). In MM-DiT blocks, parameters are divided into three components: joint text attention (Txt-Attn), joint image attention (Img-Attn), and additional modules like multi-layer perception (MLP) and modulation blocks. Figure 3 shows that MM-DiTs have approximately double the weight change of Single-DiTs, with the Txt-Attn component showing nearly twice the change of other MM-DiT elements, despite it is only 5% parameters of total backbone.
[…]
@@
To fill the gap in high-quality datasets that combine text with typographic attributes, we created the TC-Dataset using typography control rendering (TC-Render). The pipeline leverages HTML rendering to produce images featuring typographic attributes like fonts and word-level styles such as bold, italic, and underline, as shown in Figure 4. Details of the dataset are available in the supplementary Sec 5.
@@ 3.2 Style Control Adapters
$\begin{split}{z}^{\prime\prime}=\text{Attention}({Q}^{\prime},{K}^{\prime},{V}^{\prime})=\text{Softmax}(\frac{{Q^{\prime}}{K^{\prime}}^{\top}}{\sqrt{d}}){V}^{\prime},\\ \end{split}$ (4)
@@
where ${Q}={z_{c}}{W}_{q},{K}={z_{c}}{W}_{k},{V}={z_{c}}{W}_{v},{Q}^{\prime}={z_{t}}{W}_{q},{K}^{\prime}={c}_{img}{W}^{\prime}_{k},{V}^{\prime}={c}_{img}{W}^{\prime}_{v}$ , and $\lambda$ represents scale of $c_{img}$. And only ${W}^{\prime}_{k}$ and ${W}^{\prime}_{v}$ are trainable.
Style Control Training. Style control training consists of two phases, each utilizing different carefully prepared datasets. The phase 1 involves common pretraining with general image-text pairs. Besides, we have introduced phase 2 to better adapt to ATR tasks and avoid content leakage caused by using artistic text images as input. The phase 2 is further fine-tuning after phase 1, using a dataset that includes artistic text images and paired descriptions. For phase 1, we assembled a dataset called SC-general, which includes approximately 580k general image-text pairs with high aesthetic scores. These images were sourced from open-source datasets [14, 37]. For phase 2, we created the SC-artext dataset. We compile a list of style descriptions and a list of words. Combining these lists generated various prompts for artistic text images, which were then used as input for Flux.1-dev [4], resulting in approximately 20k high-quality images. To ensure the images matched the original text content, we used shareGPT4v [11] to regenerate captions. The datasets are detailed in supplementary Sec 5.
Design Choice of Image Encoder. In artistic text rendering, borrowing the style of artistic text images is crucial [44, 45, 46]. But these images carry text information, risking content leakage. To avoid this, the image encoder should be as text-agnostic as possible. Therefore, we select CLIP [33] among alternatives like DINO [30], Resampler [2] or SigLIP [50] widely used in existing adapers [12, 49], due to CLIP’s visual embeddings are text-insensitive [23, 12]. More discussion are in supplementary Sec 1.2-(3).
4 Experiments
Text Rendering Benchmark. To assess the text rendering capabilities with word-level typography and style controls, we extend the existing scene text rendering benchmark [9] by introducing new benchmarks for basic text and artistic text rendering. BTR-bench. To evaluate word-level typography controls in basic text rendering, we introduce basic text rendering benchmark (BTR-bench). BTR-bench includes 100 prompts of different fonts and typographic attributes. For each text prompt, typographic attributes are randomly applied to three positions within the text to assess the model’s ability to render specific typographic attributes on individual words, while font attributes are applied to the entire text in the image. ATR-bench. To evaluate artistic text rendering, we introduce the Artistic Text Rendering benchmark (ATR-bench). Similar to the single-letter and multi-letter classification in [38], we categorize the content into single-word and multi-word groups. Drawing on the style prompts from the GenerativeFont benchmark [28], we generate artistic individual letters and words using Flux [1]. These generated artistic letters and words are used for single-word and multi-word text rendering, respectively.
Implementation Details. We use Flux.1-dev [1] as the base model for its strong text rendering. In TC-FT, we fix the text prompt for CLIP text encoder since the pooled embedding provides only coarse-graine information [15].
Training details. For TC-FT, we fine-tune the base model for 40k steps using the TC-Dataset, incorporating a regularization prefix (‘sks’) in the text prompts. The total batch size is 32. For the style control adapters, we train for 100k steps on SC-general and 15k steps on SC-artext dataset, using total batch size of 64. All training is on 8 $\times$ A100 and 512 resolution, with a learning rate of $1\times 10^{-5}$.
Methods Consistency Accuracy
@@ 4 Experiments
Glyph [23] 93.68 32.73 96.36
TD-2 [10] 86.17 1.81 42.86
SD3 [15] 87.37 0.91 48.05
Flux [1] 90.67 0.91 66.49
Ours 96.98 63.64 55.00 82.85
Table 2: Quantitative results for basic text rendering.
4.1 Quantitative Results
Quantitative Metrics. We conduct evaluations from two perspectives: consistency and accuracy. For accuracy, we use tool [5] in [41] to calculate the OCR accuracy (OCR-Acc). In the basic text rendering (BTR), existing OCR tools struggle to evaluate word-level typographic attribute accuracy (Word-Acc). Therefore, we use GPT4o and manual screening to assess and obtain the corresponding score. For consistency, we calculate CLIP image scores (CLIP-I) in artistic text rendering (ATR) and scene text rendering (STR). To better evaluate font consistency, we use FontCLIP [39] instead of CLIP, referring to the scores as FontCLIP-I. Beyond automated evaluations, we conduct user studies for font consistency (Font-Con) in BTR and style consistency (Style-Con) in ATR.
Basic Text Rendering. We compare with Glyph-ByT5 (Glyph) [23], TextDiffuser-2 (TD-2) [10], SD3-medium (SD3) [15] and Flux.1-dev (Flux) [1] on BTR-bench. As shown in Table 2, our method outperformed the baselines in three out of four metrics while slightly below Glyph-ByT5 regarding OCR accuracy in BTR. This is reasonable since Glyph-ByT5 was trained on millions of text images, whereas our approach utilized a dataset of only 50k basic text images, which is twenty times smaller than theirs.
Methods Consistency Accuracy
@@ 4.1 Quantitative Results
SD3 [15] 60.24 13.64 24.16
Flux [1] 64.06 17.42 48.71
SD3-IPA 59.59 0 / 9.09 6.25 / 9.18
Flux-IPA 57.05 3.41 / 2.27 19.14 / 28.49
SD3-IPA+SC 62.66 20.12 / 22.43 5.40 / 14.34
Flux-IPA+SC 63.57 23.35 / 25.76 13.02 / 30.15
Flux-Redux 59.56 0 / 10.32 0 / 5.83
Ours 64.27 31.82 / 34.09 61.78 / 59.66
Table 3: Quantitative comparison of artistic text rendering. Ours, SD3-IPA, and Flux-IPA use scale = 0.9/0.6. Redux considers original / interpolation settings. For CLIP-I, the average is reported. SC: style captions. Text prompts for each method in Figure 6.
Artistic Text Rendering. Comparing with SD3 [15], Flux [1], SD3 with IP Adapter (SD3-IPA)  [^1] SD3-IPA, Flux-IPA  [^2] Flux-IPA, Flux-Redux  [^3] Flux-Redux, results are shown in Table 3. Moreover, we refer to the ComfyUI Node  [^4] ComfyUI-AdvancedRefluxControl and apply interpolation to balance the image prompt with text prompt in Flux-Redux. Despite using a simpler text input, our method outperforms others across all metrics under various hyperparameter settings.
Additionally, we compare with Flux in STR task on the MARIO-bench [9]. As indicated in Table 4, our method significantly improves OCR-Acc and CLIP scores. Upon reviewing the image results, we found that Flux exhibits semantic confusion in the STR task (on MARIO-bench [9]), which notably reduces its OCR accuracy.
[…]
Flux [1] 24.17 29.93
@@
User Studies. We conducted user studies to perceptually evaluate our results against baselines. The evaluation centered on two key aspects: font consistency (Font-Con) and style consistency (Style-Con). Details are provided in the supplementary Sec 7 .
Figure 6: Qualitative comparison of style consistency and content accuracy in artistic text rendering against baselines. For all rows except the last row, the input consists of a text prompt along with style images on the top-left. In the top three rows, the text prompts are just simple captions “Text:‘Word’”, while for others are style captions.
4.2 Qualitative Results
Basic Text Rendering. We use a set of challenging prompt words for evaluation. For Flux and Glyph-ByT5, the prompt (for the leftmost images) is: “Blue Text: ‘Love knows no limits’ in Font: Josefin Sans, Add underline to ‘Love’, Background: pure yellow”. This prompt specifies the font and applies a typographic attribute to a word. In Figure 5, Glyph-ByT5 achieves better font consistency than Flux but lacks word-level control. In contrast, our method ensures strong font consistency and enables word-level control, such as underline, bold, or italic.
Artistic Text Rendering. To ensure a fair comparison, we set the same seed for each row. The style caption is the same prompt which used to generate the artistic single letters by Flux ( ‘A’ and ‘n’ in top left of Figure 6).
Obviously, our results show the best style consistency while preserving the accuracy of the text, comparing with baselines in Table 3. In second and third rows of Figure 6, because the text prompt is relatively simple, output suffers from severe content leakage from style image. There is also semantic confusion, e.g., words ‘Parrots’ becoming parrots itself. In fourth and fifth rows, after using the style caption, the text content becomes prominent. However, the style consistency remains poor, and there are issues with content and capitalization errors, such as ‘Parots’ and ‘WINDS’). When scale is set to 0.9, content leakage still exists, e.g. the first image in the fourth row (similar to ‘A’), and the fourth image in the fifth row (similar to ‘N’). In Flux-Redux, the results from original is merely about generating variations from style images, and style of results from interpolation is obviously inconsistent. The caption only method lacks style control. As a result, even with the same seed and prompt, the outputs are also inconsistent in style. In addition, we find that the typography controls in BTR can be transferred to ATR and STR to a certain extent in Figure  1. It is reasonable that the degree of controllability will be affected by given style image, particularly when it contains text. However, considering we only use basic text images to learn those word-level attributes, it shows potential for domain generalization ability of proposed method.
Figure 7: Ablation study of style control adapter (SCA) on second phase finetuing with SC-artext (Art-FT) and TC-FT.
4.3 Ablation study
Ablation on TC-FT. To assess the effectiveness of typography control fine-tuning (TC-FT), we set up four configurations: training new tokens only, T5 with new tokens, joint text attention (Txt-Attn) with new tokens, and joint text and image attention (Txt+Img-Attn) with new tokens. The results in Table 5 show the performance in the BTR task. The result of training tokens only is similar to the original Flux. This is likely due to T5 being trained solely on the text modality, lacking the joint vision-language space of CLIP. As stated in [15, 23], text rendering capabilities mainly depend on the text encoder. So we tried to train T5, we found it severely degraded text accuracy (visual results detailed in supplementary Sec 2.2). The data in the last two rows indicate that fine-tuning only Txt-Attn is a more effective approach. In training, more parameters typically demand more training steps for better performance. The text rendering capabilities of SD3/Flux also rely on DPO (Direct Preference Optimization) [15], which we didn’t apply. Without DPO, fewer training steps are needed to preserve prior knowledge and mitigate overfitting. As shown in Figure 10, excessive training steps reduced the in-context nature of text and background in scene text images